OpenStudy (anonymous):

find an equation of the line containing the given par of points (-2,0) and (0,1)

OpenStudy (anonymous):

will someone please help me

OpenStudy (anonymous):

write slope intercept from?

OpenStudy (anonymous):

(-2,0) and (0,1)

OpenStudy (anonymous):

whats the equation of the line

OpenStudy (anonymous):

m=(1-0)/(0-(-2) m=1/2 y-(-2)= 1/2(x-0)\[y=\frac{1}{2}x+2\]

OpenStudy (anonymous):

is that the final answerr

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thnx